Are the following lines parallel, perpendicular, or neither?
L1 with equation x-2y=10
L2 with equation 2x+y=2

Answer by stanbon(75887)

Are the following lines parallel, perpendicular, or neither?
L1 with equation x-2y=10
L2 with equation 2x+y=2
---------
Write in slope-intercept form, as follows:
L1: y=(1/2)x-5
L2: y=-2x+2
Since the slopes are -1/2 and 2 the lines are perpendicular.
Cheers,
Stan H.